Project

General

Profile

11082013

Nightly Observing Plan for MJD 56605 (starting evening of Nov 08, 2013)

Observers

Observer 1 Observer 2 Run Manager SISPI Expert on Call
Yuanyuan Zhang Liz Buckley-Geer Marisa March

CTIO staff (Telescope Operators)

Day Support Night Assistant
Alberto Alvarez Leonardo Paredes

DECam Configuration and Setup

SISPI Configuration File DES.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0001
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac
mjd 56605, date 20131108, Site CTIO, UT-offset 3 hr
Current local time 14:56 CLST     UT 17:56     SidTim 16:26
Sunset     20:19 CLST  23:19 UT   SidTim  327 deg  
Eve. -10   20:57 CLST  23:57 UT   SidTim  337 deg  Zeropoint; standards
Eve. -14   21:18 CLST  00:18 UT   SidTim  342 deg  Begin obstac
Eve. -18   21:39 CLST  00:39 UT   SidTim  347 deg  Last chance standards
Moonset    01:18 CLST  04:18 UT   SidTim   42 deg  
Midpoint   01:27 CLST  04:27 UT   SidTim   44 deg  
Morn -18   05:14 CLST  08:14 UT   SidTim  101 deg  Begin astronomical twilight
Morn Std   05:35 CLST  08:35 UT   SidTim  106 deg  End obstac; begin standards
Morn -10   05:56 CLST  08:56 UT   SidTim  112 deg  
Sunrise    06:34 CLST  09:34 UT   SidTim  121 deg  
Moon position (deg): RA 303.1, Dec -13.9 Phase: 0.425

Weather Conditions

*Tololo Weather Links .

http://www.meteochile.gob.cl/cna/irsat_alta.png

Observing Plan and Goals for the Night

  1. Afternoon calibrations script to use: prenight_calibration.json
  2. Start of night standards
  3. ObsTac
  4. End of night standards
  5. End of night calibrations

Start of Night Standards
UT Date = '2013/11/08 23:57' Local time = 20:57
SDSSJ0200-0000 ra 02:00:00 dec +00:00:00 secz 1.9 script to use DES_std_sdssj0200_Yzirg.json
SDSSJ0100-0000 ra 01:00:00 dec +00:00:00 secz 1.5 script to use DES_std_sdssj0100_Yzirg.json
SDSSJ2300-0000 ra 23:00:00 dec +00:00:00 secz 1.2 script to use DES_std_sdssj2300_Yzirg.json

End of Night Standards
UT Date = '2013/11/09 08:35' Local time = 05:35
SDSSJ0320-0000 ra 03:20:00 dec +00:00:00 secz 2.1 script to use DES_std_sdssj0320_grizY.json
SDSSJ0958-0010 ra 09:58:00 dec -00:10:00 secz 1.6 script to use DES_std_sdssj0958_grizY.json
MaxVis ra 06:30:00 dec -58:45:00 secz 1.1 script to use DES_std_maxvis_grizY.json

Notes from the 4PM Meeting and Additional Comments:

Run Manager's Meeting 11/08/2013
--------------------------------
Alberto, Manuel, Marisa, Eric, Yuanyuan, Liz, Steve, Douglas, Tom
Josh (lurking), Robert, Michael, Margaret

Telescope and Camera Status (Alberto)
-------------------------------------
Everything OK. Ready to start dome flats. During the day they
have been doing some pumping on the LN2 vacuum lines, with significant improvement
at the cryocoolers.

The M and C-floor cooling hasn't been operating at full efficiency this season. That is being
improved. One temperature sensor, previously temporarily located on the C-floor,
was relocated to FNAL in mid-October. That analog sensor is noisy at times.
Digital Temperature sensors are being moved to the cooling coils themselves.

Last Night's Data (Marisa and Liz)
----------------------------------
Last night backplane 5 didn't readout at the start during bias frames. So
that's a PanD role restart. Maybe we need Telops take a few vsub off bias frames
to check this. Liz put details in the elog.

One stuck shutter error required a reset and a reconfigure.

Lots of decent WF data last night.

Every 5 or 10 minutes we get a warning beep about more than 7 degrees separation.
Maybe check domeaz vs telescope az after a long slew. Kentools can pull up the
donuts.

DESDM (Margaret)
----------------
There was a problem ingesting the files from last night into the database.
One file is still missing. Greg is looking into that.

11/5 and 11/6 are both close to being done with DQ ready tomorrow.
That should be very interesting.

Tonight
-------------------
Obstac has two tweaks:
EN noticed that we've been doing a lot of tiling 2.
It hasn't been configured to do Y when the seeing was good. A few sections
of tiling 2 Y are close to getting near setting. So it goes to 2 but then
because the seeing is good it hasn't done the Y. That needs to be fixed right now.
This is the first time the weather has been good-enough to notice this bug.
Then When the Moon either rises or sets Obstac changes filters. Eric wants Obstac
to be able to go to be able to change to a new tiling.
The new code is checked-in and needs to be tagged and installed, then a SISPI restart.
Eric will declare it while the calibs are going on and SISPI will get a restart
when observers get back from dinner.

Weather looks good right now. Clouds in the S, as was so last night.
Eric expects Obstac will start with tiling 2 and then move to tiling 3 when the moon sets.
SK reports he looks at
http://rammb.cira.colostate.edu/ramsdis/online/images/latest/rmtc/rmtcsasec1ir404.gif

AOB
---
4 PM Meeting is on for tomorrow.